# Zihan Zhang Princeton

### v4

@@ B.3 Proof of Lemma 3
Set $\delta^{\prime}=\delta/(32T^{4}k^{2})$, and let
Let us define
$\overline{\mathcal{W}}$ $\coloneqq[k]/\cup_{j}\mathcal{W}_{j}.$ (68b)
Lemma 13.
holds all $1\leq j\leq\overline{j}$, with $\overline{j}$ defined in (67).
Importantly, this lemma tells us that, with probability at least $41-8T^{4}k^{2}\delta^{\prime}=1-\delta/$ 4, one has
Noteworthily, proving Lemma 13 is the most challenging part of our analysis, and we dedicate the next section (Appendix C) to the proof of Lemma 13.
Appendix C Controlling the Hedge trajectory (proof of Lemma 13)
This section is devoted to proving Lemma 13. The proof relies heavily on the concepts of “segments” introduced in Section 4.2. For convenience, we restate the definition below.
Definition 4 (Segment (restated)).
  • $\mathrm{(i)}$
    $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[p/2,p\right]$,
  • $\mathrm{(ii)}$
    $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$,
  • $\mathrm{(iii)}$
    $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$.
We shall refer to $t_{1}$ as the starting point and $t_{2}$ as the end point, and call $\mathcal{I}$ the index set. Moreover, two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are said to be disjoint if either $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$.
In addition, throughout this section, we shall take $\delta^{\prime}=\delta/(32T^{4}k^{2})$, and focus on any $j$ obeying
$2\displaystyle 1\leq j\leq\log_{2}(k)-$ 2 . (70)
C.1 Main steps of the proof
In this subsection, we present the main steps of the proof. Before continuing, we find it helpful to underscore a high-level idea: if $|\mathcal{W}_{j}|$ is large, then there exist many disjoint segments, thereby requiring the total length $T$ to be large enough in order to contain these segments. The key steps to construct such disjoint segments of interest are as follows:
  1. 1.
    Construct a suitable segment for each $i\in\mathcal{W}_{j}$ (see Lemma 14);
  2. 2.
  3. 3.
C.1.1 Step 1: showing existence of a segment for each distribution in $\mathcal{W}_{j}$
Lemma 14.
Proof.
$\tau=\max\big{\{}t\mid t\leq e_{i},w^{t}_{i}\leq 2^{-(j+2)}\big{\}},$
C.1.2 Step 2: constructing disjoint segments with good coverage
Lemma 15.
  1. $\mathrm{(i)}$
  2. $\mathrm{(ii)}$
  3. $\mathrm{(iii)}$
Proof.
See Appendix C.2. ∎
Lemma 16.
Recall the definition of $\mathcal{W}_{j}$ in (68a). For each $i\in\mathcal{W}_{j}$, denote by $(s_{i},e_{i})$ the segment identified in Lemma 14. Then there exists a group of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ of $\mathcal{W}_{j}$ satisfying the following properties:
  1. $\mathrm{(i)}$
  2. $\mathrm{(ii)}$
  3. $\mathrm{(iii)}$
    • $\mathrm{(a)}$
    • $\mathrm{(b)}$
    • $\mathrm{(c)}$
Proof.
See Appendix C.3. ∎
In brief, the subsets $\{\mathcal{V}_{j}^{n}\}$ identified in Lemma 16 enjoy the following useful properties: (a) they are disjoint; (b) these subsets taken collectively also cover a reasonably large fraction of the elements in $\mathcal{W}_{j}$; (c) each subset $\mathcal{V}_{j}^{n}$ is linked with a suitable segment, whose associated weights have increased sufficiently from its starting point to its end point.
C.1.3 Step 3: bounding the length of segments
Recall the definition (58) of $v^{t}$ as follows
$v^{t}\coloneqq L(h^{t},w^{t})-\mathsf{OPT}\qquad\text{with }\mathsf{OPT}\coloneqq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i}),$ (73)
Lemma 17.
$t_{2}-t_{1}\geq\frac{x}{2\eta}.$ (74)
Moreover, if
$\frac{qx^{2}}{50\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\geq\frac{1}{k}$ (75)
  • (a)
    the length of the segment satisfies
  • (b)
    the quantities $\{v^{t}\}$ obey
Proof.
See Appendix C.4. ∎
C.1.4 Step 4: putting all this together
To finish up, it follows from Property (ii) of Lemma 16 that
C.2 Proof of Lemma 15
For any integer $11\leq x\leq\log_{2}(T)+$ 1, define
@@ C.2 Proof of Lemma 15
so that the length of each segment associated with $\mathcal{W}_{j}(x)$ lies within $[2^{x-1},2^{x}]$. Let $x^{\star}$ indicate the one that maximizes the cardinality of $\mathcal{W}_{j}(x)$:
@@ C.3 Proof of Lemma 16
In what follows, we take [^8] We assume $\widetilde{Q}\geq 2$ without loss of generality. In the case $\widetilde{Q}=1$, we simply choose an arbitrary element in $\widehat{\mathcal{W}}_{j}^{p}$ as a single subset. In this way, we can collect at least $\frac{1}{4}|\widehat{\mathcal{W}_{j}^{p}}|$ segments.
@@
Step b): constructing $\mathcal{K}_{l}$ and $\widetilde{\mathcal{W}}_{j}^{p}(l)$. Let us take $e_{0}=t_{\mathrm{mid}}^{p}$, and employ $[e_{0},e_{k}]\oplus a$ as a shorthand notation for $[e_{a},e_{k+a}]$. We can then define a group of disjoint intervals of $[T]$ as follows:
$\mathcal{K}_{1}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-1}}]\right\};$ (93a)
$\mathcal{K}_{2}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-2}}],[e_{0},e_{2^{\widetilde{l}-2}}]\oplus 2^{\widetilde{l}-1}\right\};$ (93b)
$\mathcal{K}_{3}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-3}}],[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 2^{\widetilde{l}-2},[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 2\cdot 2^{\widetilde{l}-2},[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 3\cdot 2^{\widetilde{l}-2}\right\};$ (93c)
@@
$\mathcal{K}_{l}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-l}}],[e_{0},e_{2^{\widetilde{l}-l}}]\oplus 2^{\widetilde{l}-l+1},[e_{0},e_{2^{\widetilde{l}-l}}]\oplus 2\cdot 2^{\widetilde{l}-l+1},\dots,[e_{0},e_{2^{\widetilde{l}-l}}]\oplus(2^{l-1}-1)2^{\widetilde{l}-l+1}\right\};$ (93d)
@@
$\mathcal{K}_{\widetilde{l}}$ =$=\left\{[e_{2i},e_{2i+1}]\mid i=0,1,2,\dots,2^{\widetilde{l}-1}-1\right\};$ (93e)
$\mathcal{K}_{\widetilde{l}+1}$ =$=\left\{[e_{2i+1},e_{2i+2}]\mid i=0,1,2,\dots,2^{\widetilde{l}-1}-1\right\}.$ (93f)
@@
$[e_{\mathsf{trunc}(i,l^{\star}-1)},e_{\mathsf{trunc}(i,l^{\star})}]\in\mathcal{K}_{l^{\star}},$
@@
In words, $[\widehat{s}(p,o),\widehat{e}(p,o)]$ can be understood as the $o$-th interval in the set $\mathcal{K}_{l^{\star}(p)}$.
@@
To finish up, it remains to verify that the intersection of $[\widehat{s}(p_{1},o_{1}),\widehat{e}(p_{1},o_{1})]$ and $[\widehat{s}(p_{2},o_{2}),\widehat{e}(p_{2},o_{2})]$ is either empty or contains only the boundary points, unless $(p_{1},o_{1})=(p_{2},o_{2})$. To show this, note that in the case where $p_{1}\neq p_{2}$ (assuming $p_{1}<p_{2}$), we have
@@ C.4 Proof of Lemma 17
We start by proving the first claim (74). Recall that $[t_{1},t_{2}]$ is assumed to be a $(p,q,x)$-segment. From the definition of the segment (see Definition 4), there exists $i\in[k]$ such that
@@
=$\overset{\mathrm{(i)}}{=}\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w_{i}^{\tau}\exp(\eta\widehat{r}_{i}^{\tau})\right)\overset{\mathrm{(ii)}}{\leq}\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w_{i}^{\tau}+\sum_{i=1}^{k}w_{i}^{\tau}(\eta\widehat{r}_{i}^{\tau})+2\sum_{i=1}^{k}w_{i}^{\tau}\eta^{2}(\widehat{r}_{i}^{\tau})^{2}\right)$
@@
=$\overset{\mathrm{(iv)}}{=}\eta\sum_{\tau=t}^{t_{2}-1}v^{\tau}+\eta(t_{2}-t)\mathsf{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}\big{\langle}w_{i}^{\tau},\widehat{r}_{i}^{\tau}-v_{h^{\tau},i}-\mathsf{OPT}\big{\rangle}+2(t_{2}-t)\eta^{2}$
@@
By definition, we know that $\tau_{1}=t_{1}$ and $\tau_{j_{2}}\geq\tau_{j_{1}}$ for $j_{2}\geq j_{1}$. Let $\mathcal{I}$ be the index set associated with this segment $[t_{1},t_{2}]$, and set $y_{j}\coloneqq\sum_{i\in\mathcal{I}}w_{i}^{\tau_{j}}$. We then claim that there exists $1\leq\widetilde{j}\leq j_{\mathrm{max}}$ such that
$\log\left(\frac{y_{\widetilde{j}+1}}{y_{\widetilde{j}}}\right)$ $\geq\frac{x}{\log_{2}(1/\eta)+1}.$ (116)
@@
  • Case 2: $11\leq\widetilde{j}\leq j_{\mathrm{max}}-$ 1. It comes from the definition (115) that
@@
    When $11\leq\widetilde{j}\leq j_{\mathrm{max}}-$ 1, the above display taken collectively with (120) gives
@@
C.5 Additional illustrative figures for segment construction
@@
  • In Figure 5 and Figure 6, we illustrate how to construct disjoint segments from a group of segments with common starting points in the case with $k=8$. In this particular example, we have in total $5$ groups of disjoint segments, marked with different colors.
@@
Appendix D Proofs of the lower bound in Theorem 2
D.1 Proof of Theorem 2
@@ D.2 Proof of Lemma 18
where (129) results from Lemma 19 in Appendix D.3, and (128) and (130) hold since for $h^{\prime}=h,h^{\star}$, the event $\{h_{\mathrm{out}}=h^{\prime},M_{\widetilde{i}}=m^{\prime}\}$ is independent of $\{x^{l}(\widetilde{i})\}_{l\geq m^{\prime}+1}$. Taking the sum over $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we obtain
@@
D.3 Statement and proof of Lemma 19
@@
Consider a given sequence $\{m_{i}\}_{i=1}^{k}$. Let $X(i)=\{X^{l}(i)\}_{l=1}^{m_{i}}\in\{-1,0,1\}^{kNm_{i}}$ for $i\in[k]$, and let $x(i)=\{x^{l}(i)\}_{l=1}^{m_{i}}$ be the datapoints of the first $m_{i}$ calls to $\mathsf{Query}(D_{i})$. With slight abuse of notation, we take $\sigma(x(i))=\{\sigma(x^{l}(i))\}_{l=1}^{m_{i}}$ for each $i\in[k]$. It then follows from Lemma 20 in Appendix D.4 that
@@
D.4 Statement and proof of Lemma 20
@@
Appendix E Proofs of auxiliary lemmas for Rademacher classes
E.1 Proof of Lemma 6
@@
~=$\widetilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\},$
@@
for each $1\leq t\leq T$. Taking $\delta^{\prime}=\delta/4$, we obtain that with probability at least $41-\delta/$ 4,
@@ E.2 Proof of Lemma 4
=$\qquad\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
@@ E.3 Proof of Lemma 5
with probability exceeding $1-1/(2k)$. Defining $\mathcal{E}$ to be the event that $\widehat{n}_{i}\geq n_{i}/6$ holds for all $i\in[k]$, we can invoke the union bound to see that
@@ E.4 Necessity of Assumption 1
=$\mathcal{H}=\big{\{}\mathsf{joint}(\{h_{i}\}_{i=1}^{k})\mid h_{i}\in\mathcal{H}_{i},\forall i\in[k]\big{\}}.$ (146)
@@
$\mathbb{P}_{\mathcal{D}^{\prime}_{n^{*}}}\{x_{n^{*}}\}$ $=\frac{1}{2}\mathds{1}\{x_{n^{*}}=1\}+\frac{1}{2}\mathds{1}\{x_{n^{*}}=0\}$ for some $\text{for some }n^{*};$ (152)
@@
$\mathbb{P}_{\mathcal{D}^{\prime}_{n}}\{x_{n}\}$ $=\left(\frac{1}{2}+2\varepsilon\right)\mathds{1}\{x_{n^{*}}=1\}+\left(\frac{1}{2}-2\varepsilon\right)\mathds{1}\{x_{n^{*}}=0\}$ for all $\text{for all }n\neq n^{*}.$ (153)

### v5

@@ B.3 Proof of Lemma 6
For any integer $1\leq x\leq\log_{2}(T)+1$, define
@@
so that the length of each segment associated with $\mathcal{W}_{j}(x)$ lies within $2^{x-1},2^{x}]$. Let $x^{\star}$ indicate the one that maximizes the cardinality of $\mathcal{W}_{j}(x)$:
@@ B.4 Proof of Lemma 7
[…]
@@
Step b): constructing $\mathcal{K}_{l}$ and $\widetilde{\mathcal{W}}_{j}^{p}(l)$. Let us take $e_{0}=t_{\mathrm{mid}}^{p}$, and employ $e_{0},e_{k}]\oplus a$ as a shorthand notation for $e_{a},e_{k+a}]$. We can then define a group of disjoint intervals of $[T]$ as follows:
$\mathcal{K}_{1}$ $=\left\{[e_{0},e_{2^{\widetilde{l}-1}}]\right\};$ (98a)
$\mathcal{K}_{2}$ $=\left\{[e_{0},e_{2^{\widetilde{l}-2}}],[e_{0},e_{2^{\widetilde{l}-2}}]\oplus 2^{\widetilde{l}-1}\right\};$ (98b)
$\mathcal{K}_{3}$ $=\left\{[e_{0},e_{2^{\widetilde{l}-3}}],[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 2^{\widetilde{l}-2},[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 2\cdot 2^{\widetilde{l}-2},[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 3\cdot 2^{\widetilde{l}-2}\right\};$ (98c)
@@
$\mathcal{K}_{l}$ $=\left\{[e_{0},e_{2^{\widetilde{l}-l}}],[e_{0},e_{2^{\widetilde{l}-l}}]\oplus 2^{\widetilde{l}-l+1},[e_{0},e_{2^{\widetilde{l}-l}}]\oplus 2\cdot 2^{\widetilde{l}-l+1},\dots,[e_{0},e_{2^{\widetilde{l}-l}}]\oplus(2^{l-1}-1)2^{\widetilde{l}-l+1}\right\};$ (98d)
@@
$\mathcal{K}_{\widetilde{l}}$ $=\left\{[e_{2i},e_{2i+1}]\mid i=0,1,2,\dots,2^{\widetilde{l}-1}-1\right\};$ (98e)
$\mathcal{K}_{\widetilde{l}+1}$ $=\left\{[e_{2i+1},e_{2i+2}]\mid i=0,1,2,\dots,2^{\widetilde{l}-1}-1\right\}.$ (98f)
@@
$e_{\mathsf{trunc}(i,l^{\star}-1)},e_{\mathsf{trunc}(i,l^{\star})}]\in\mathcal{K}_{l^{\star}},$
@@
In words, $\widehat{s}(p,o),\widehat{e}(p,o)]$ can be understood as the $o$-th interval in the set $\mathcal{K}_{l^{\star}(p)}$.
@@
To finish up, it remains to verify that the intersection of $\widehat{s}(p_{1},o_{1}),\widehat{e}(p_{1},o_{1})]$ and $\widehat{s}(p_{2},o_{2}),\widehat{e}(p_{2},o_{2})]$ is either empty or contains only the boundary points, unless $(p_{1},o_{1})=(p_{2},o_{2})$. To show this, note that in the case where $p_{1}\neq p_{2}$ (assuming $p_{1}<p_{2}$), we have
@@ B.5 Proof of Lemma 8
We start by proving the first claim (39). Recall that $t_{1},t_{2}]$ is assumed to be a $(p,q,x)$-segment. From the definition of the segment (see Definition 1), there exists $i\in[k]$ such that
@@
$\overset{\mathrm{(i)}}{=}\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w_{i}^{\tau}\exp(\eta\widehat{r}_{i}^{\tau})\right)\overset{\mathrm{(ii)}}{\leq}\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w_{i}^{\tau}+\sum_{i=1}^{k}w_{i}^{\tau}(\eta\widehat{r}_{i}^{\tau})+2\sum_{i=1}^{k}w_{i}^{\tau}\eta^{2}(\widehat{r}_{i}^{\tau})^{2}\right)$
@@
$\overset{\mathrm{(iv)}}{=}\eta\sum_{\tau=t}^{t_{2}-1}v^{\tau}+\eta(t_{2}-t)\mathsf{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}\big{\langle}w_{i}^{\tau},\widehat{r}_{i}^{\tau}-v_{h^{\tau},i}-\mathsf{OPT}\big{\rangle}+2(t_{2}-t)\eta^{2}$
@@
By definition, we know that $\tau_{1}=t_{1}$ and $\tau_{j_{2}}\geq\tau_{j_{1}}$ for $j_{2}\geq j_{1}$. Let $\mathcal{I}$ be the index set associated with this segment $t_{1},t_{2}]$, and set $y_{j}\coloneqq\sum_{i\in\mathcal{I}}w_{i}^{\tau_{j}}$. We then claim that there exists $1\leq\widetilde{j}\leq j_{\mathrm{max}}$ such that
[…]
@@
  • Case 2: $1\leq\widetilde{j}\leq j_{\mathrm{max}}-1$. It comes from the definition (120) that
@@
    When $1\leq\widetilde{j}\leq j_{\mathrm{max}}-1$, the above display taken collectively with (125) gives
@@
B.6 Additional illustrative figures for segment construction
@@
  • In Figure 5 and Figure 6, we illustrate how to construct disjoint segments from a group of segments with common starting points in the case with $k=8$. In this particular example, we have in total 55 groups of disjoint segments, marked with different colors.
@@
Appendix C Proofs of the lower bound in Theorem 2
C.1 Proof of Theorem 2
@@ C.2 Proof of Lemma 18
where (134) results from Lemma 19 in Appendix C.3, and (133) and (135) hold since for $h^{\prime}=h,h^{\star}$, the event $\{h_{\mathrm{out}}=h^{\prime},M_{\widetilde{i}}=m^{\prime}\}$ is independent of $\{x^{l}(\widetilde{i})\}_{l\geq m^{\prime}+1}$. Taking the sum over $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we obtain
@@
C.3 Statement and proof of Lemma 19
@@
Consider a given sequence $\{m_{i}\}_{i=1}^{k}$. Let $X(i)=\{X^{l}(i)\}_{l=1}^{m_{i}}\in\{-1,0,1\}^{kNm_{i}}$ for $i\in[k]$, and let $x(i)=\{x^{l}(i)\}_{l=1}^{m_{i}}$ be the datapoints of the first $m_{i}$ calls to $\mathsf{Query}(D_{i})$. With slight abuse of notation, we take $\sigma(x(i))=\{\sigma(x^{l}(i))\}_{l=1}^{m_{i}}$ for each $i\in[k]$. It then follows from Lemma 20 in Appendix C.4 that
@@
C.4 Statement and proof of Lemma 20
@@
Appendix D Proofs of auxiliary lemmas for Rademacher classes
D.1 Proof of Lemma 11
@@
$\widetilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\},$
@@
for each $1\leq t\leq T$. Taking $\delta^{\prime}=\delta/4$, we obtain that with probability at least $1-\delta/4$,
@@ D.2 Proof of Lemma 9
$\qquad\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
@@ D.3 Proof of Lemma 10
with probability exceeding 11/(2k)1-1/(2k). Defining $\mathcal{E}$ to be the event that $\widehat{n}_{i}\geq n_{i}/6$ holds for all $i\in[k]$, we can invoke the union bound to see that
@@ D.4 Necessity of Assumption 1
$\mathcal{H}=\big{\{}\mathsf{joint}(\{h_{i}\}_{i=1}^{k})\mid h_{i}\in\mathcal{H}_{i},\forall i\in[k]\big{\}}.$ (151)
@@
$\mathbb{P}_{\mathcal{D}^{\prime}_{n^{*}}}\{x_{n^{*}}\}$ $=\frac{1}{2}\mathds{1}\{x_{n^{*}}=1\}+\frac{1}{2}\mathds{1}\{x_{n^{*}}=0\}$ $\text{for some }n^{*};$ (157)
@@
$\mathbb{P}_{\mathcal{D}^{\prime}_{n}}\{x_{n}\}$ $=\left(\frac{1}{2}+2\varepsilon\right)\mathds{1}\{x_{n^{*}}=1\}+\left(\frac{1}{2}-2\varepsilon\right)\mathds{1}\{x_{n^{*}}=0\}$ $\text{for all }n\neq n^{*}.$ (158)